OpenStudy (anonymous):

Richard works at an icecream shop, regular cones get two scoops of icecream; large gets three scoops of icecream. Richard scoops 234 regular cones and 156 large cones. one scoop of icecream is 3 ounces. a tub of icecream weighs 10 pounds. How many tubs of icream did richard use?

OpenStudy (anonymous):

Let's see: 234 regular cones, 3 ounces a dip, 2 dips each=234 *3*2 =1404 ounces. 156 large cones 3 ounces a dip, 3 dips each: 156*3*3 =1404 ounces. 1404 +1404 =2808 ounces. !0 pounds =160 ounces; 2808/160=17.55; round up, it's 18 tubs.your answer is 18 tubs

OpenStudy (anonymous):

hope it helped

OpenStudy (anonymous):

???